# Interpreting GNN-based IDS Detections Using Provenance Graph Structural Features

### v5

@@ Abstract
Advanced cyber threats (e.g., Fileless Malware  and Advanced Persistent Threat (APT)) have driven the adoption of provenance-based security solutions. These solutions employ Machine Learning (ML) models for behavioral modeling and critical security tasks such as malware and anomaly detection. However, the opacity of ML-based security models limits their broader adoption, as the lack of transparency in their decision-making processes restricts explainability and verifiability. We tailored our solution towards Graph Neural Network (GNN)-based security solutions since recent studies employ GNNs to comprehensively digest system provenance graphs for security critical tasks.
@@
To enhance the explainability of GNN-based security models, we introduce ProvExplainer, a framework offering instance-level security-aware explanations using a interpretable surrogate model. ProvExplainer’s interpretable feature space consist of discriminant subgraph patterns and graph structural features which can be directly mapped to the system provenance problem space, making the explanations human understandable. Considering both the subgraph patterns and graph structural features, gives ProvExplainer the unique advantages of providing explanations that are sensitive to both local as well as global changes.
@@
By considering prominent GNN architectures (e.g., GAT  and GraphSAGE) for program classification and anomaly detection tasks, we show how ProvExplainer synergizes with current state-of-the-art (SOTA) GNN explainers to deliver domain and instance specific explanations. We measure the explanation quality using $\text{fidelity}^{+}$ /$\text{fidelity}^{-}$  metric as used by traditional GNN explanation literature, and we incorporate the precision/recall metric where we consider the accuracy of the explanation against the ground-truth. On malware and APT datasets, ProvExplainer achieves up to 29% / 27% / 25% higher $\text{fidelity}^{+}$ , precision and recall, and 12% lower $\text{fidelity}^{-}$  respectively, compared to SOTA GNN explainers.
1 Introduction
@@
Given the significant threat posed by advanced and sophisticated adversaries [apt0, apt1, sony, solarwinds] (i.e., malware writers and APT actors), security is paramount in today’s highly digitized society. Recent advances in system monitoring have generated a variety of fine-grained telemetry for advanced security analysis. To leverage this telemetry to counter security threats, various learning-based security tools have been proposed and deployed [dahl2013large, arp2014drebin, gandotra2014malware, saxe2015deep, grosse2016adversarial, narayanan2016performance, kolosnjaji2016deep, javaid2016deep, tang2016deep, holmes2019sp, wang2020ndss, han2020ndss, morse2020sp, sigl2021sec, cheng2024kairos, rehman2024flash, goyal2024rcaid]. System provenance captures causally dependent system events on an end host, represented as a provenance graph. Provenance graphs are heterogeneous, comprising various node types (e.g., process, file, and socket) and edge types (e.g., read, write, and send), each with textual and numerical attributes, effectively representing dynamic runtime behaviors. Hence, provenance graphs provide irreplaceable security defense for countering sophisticated and stealthy APT-like attack campaigns. Particularly, Endpoint Detection and Response (EDR) solutions [xdr0, xdr1, edr0], built on fine-grained system provenance data, have become a mainstream security defense for enterprise networks [edr1, edr2].
@@
The increasing use of provenance graphs has spurred interest in leveraging Graph Neural Networks (GNNs) for security applications. Recent advancements in GNNs enable the direct learning of relational dependencies and topological structures from graphs. GNNs utilize a permutation-invariant message-passing mechanism to aggregate information from node neighborhoods, expanding the information propagated through each neighborhood with every iteration. This allows GNNs to learn from complex graph structures automatically, which has proven beneficial for provenance-based security solutions [goyal2024rcaid, rehman2024flash, cheng2024kairos, jia2024magic]. However, while GNN-based security models are highly effective, their opaque nature significantly limits their advantages, particularly due to the lack of clear, verifiable explanations for their predictions. These models fail to provide ground-truth relevant explanations, which are essential for building trust among security practitioners. In this context, ground-truth relevant explanations refer to explanations aligned with specific system actions that represent techniques, procedures, and system artifacts that are familiar to security professionals and can be validated by existing ground-truth intelligence.
While studies [ying2019gnnexplainer, luo2020parameterized, yuan2021explainability] have explored domain-agnostic explanations for GNN decisions, they lack validation in the security domain, where contextualizing explanations with program behaviors is critical for trust in the underlying model. Recent research [warnecke2019don, warnecke2020evaluating, ganz2023hunting] has emphasized verifiable explanations for neural networks in binary analysis and vulnerability discovery, yet our study is the first to validate explanations within the system provenance domain. Traditional explainable Machine Learning (ML) techniques [xai, rudin2019stop] have focused on explanations based on the model’s input or internal parameters; however, such explanations may be inadequate or unintelligible for end-users in cybersecurity, where a single error can have catastrophic outcomes. Explanations have different meaning for model developers and end-user, since end-users cannot make sense of explanations in terms of model weights or gradients, but can make sense if explanations are in terms of edges or subgraphs. In security domain, models must explain their decision-making processes [lipton2018mythos] that are verifiable using the ground-truth.
@@
In this paper, we introduce a novel approach to enhance the explainability of GNN models based on system provenance data by integrating an interpretable surrogate model, specifically Decision Trees (DT), with security-aware discriminant subgraph patterns and graph structural features. We efficiently mined subgraph patterns [jiang2013survey] using a graph evolution miner (GERM [jiang2013survey]) to identify discriminative subgraph patterns between malicious and benign datasets. Additionally, we extracted structural graph features, such as degree centrality, betweenness centrality, closeness centrality, eigenvector centrality, clustering coefficient, and triangle clustering. These structural features capture long-term dependencies within the system, while the subgraph patterns represent local dependencies. Together, they provide explanations for both short-term and long-term dependencies in the graph. Our analysis revealed that these subgraph patterns correspond to well-known malicious activities, such as drive-by-downloads (e.g., dropper malware)[dropper, dropper2], malware staging and propagation[scar], system probing [mimikatz], and malware replication using templates [drakon]. By incorporating these features alongside graph structural measures, our model facilitates interpretability for security experts [intriguing2020sp] by closely aligning with security-relevant system behaviors.
While our research takes full advantage of an extensive dataset that we collected from real systems, we specifically focus on the structural features of system provenance graphs, without considering the text or numerical node and edge attributes. We restrict the scope of our attribute view for the following reasons: first, the core merit of provenance based Intrusion Detection System (IDS) lies in the ability to retain features drawn from causal dependencies captured in graph structural relations; these structural features are significantly more difficult for adversaries to manipulate than textual attributes (e.g., filenames). Secondly, if we cannot adequately explain GNN decisions based solely on structural features, it becomes even more challenging to interpret decisions when GNNs incorporate both structural and attribute-based complexities.
@@
Our research effort decouples the feature engineering and decision-making processes of GNN models, focusing on feature engineering to support human interpretability rather than solely improving decision accuracy. The core insight behind our study is that general-purpose explanation techniques lack security-specific considerations but can be enhanced with tailored structural features to address this gap. We emphasize that ProvExplainer is an explanation framework, not an attack detection technique. Unlike GNNs, surrogate Decision Trees (DTs) sacrifice generalizability, expressiveness, and automatic feature extraction to gain interpretability, making them unsuitable for direct attack detection. ProvExplainer is a GNN-structure-agnostic framework capable of incorporating an arbitrary number of subgraph patterns and structural graph features, and it is designed for easy extensibility with additional security-aware graph features.
We measured the effectiveness of ProvExplainer using $\text{fidelity}^{+}$  and $\text{fidelity}^{-}$  similar to previous GNN explanability research [yuan2021explainability]. We also incorporated two additional measure known as precision and recall that captures the accuracy of the explanation provided compared against the ground-truth. While ProvExplainer independently surpasses current State-Of-The-Art (SOTA) GNN explainers in the majority of our security-oriented evaluation tasks, as we show in our experimental evaluation, it can also be seamlessly integrated with other SOTA GNN explainers to develop even more effective explanations. Notably, our research does not attempt to completely replace existing GNN model explanation approaches. While stable and accountable,ProvExplainer ’s coverage is inherently limited by the capacity of our surrogate DT architecture. We intend ProvExplainer to be complementary to a large body of existing literature on GNN model explanations with specific emphasis on security domain constraints and their interpretability.
@@
To demonstrate the advantages of ProvExplainer in the system provenance domain, we compared ProvExplainer against SOTA GNN explainers such as GNNExplainer [ying2019gnnexplainer], PGExplainer [luo2020parameterized], and SubgraphX [yuan2021explainability] across well-balanced security domain tasks, including malware detection, and APT detection. These SOTA GNN explainers provide the best performance according to other security studies [ganz2021explaining, warnecke2020evaluating]. The GNN models in our evaluation are trained with an extensive dataset collected using our in-house deployment which monitored system events on an average of 20 mixed Windows and Linux hosts over 13 months. In addition, in our evaluation, we used the DARPA [darpa] Transparent Computing dataset, the APT dataset from [mukherjee2023sec], and malware lists from [survivialism2021sp].
@@
To our knowledge, our research is the first to leverage security-aware discriminant subgraph pattern and graph structural features specifically tailored for human interpretation. Broadly, we summarize the contributions of ProvExplainer as follows:
@@
  • ProvExplainer examines discriminant subgraph patterns and graph structural features to provide GNN detection explanation.
@@
  • Our security-aware features enabled surrogate DTs to achieve 95% agreement with the GNN model on malware and APT attack detection.
  • ProvExplainer improves $\text{fidelity}^{+}$  by 29%, precision by 27%, recall by 25% and decreased $\text{fidelity}^{-}$  by 12% compared to SOTA GNN explainers: GNNExplainer [ying2019gnnexplainer], PGExplainer [luo2020parameterized], and SubgraphX [yuan2021explainability].
@@
  • We curated an extensive dataset using in-house data collection, APT datasets from different sources (i.e., DARPA [darpa] and past literature [mukherjee2023sec]) and collected real-world malware samples from  [survivialism2021sp] to confirm the generalizability of ProvExplainer.
To benefit the community and facilitate future research, we are committed to publish our dataset and code.
2 Background and Related Work
We introduce system provenance and ML security applications using it. We discuss the efforts made to explain GNN models and their challenges. We discuss the verification of explanations in the system provenance domain.
Table I: Node and edge types for in-house system provenance with associated attributes.
Types Attributes
Nodes (resources) process signature, executable name, pid
file owner (uid, gid), name, inode
socket (IP) dstip, srcip, dstport, srcport, type
Edges (events) process $\rightarrow$ process command args, starttime
process $\rightarrow$ file read, write, amount
process $\rightarrow$ IP address send, recv, amount
2.1 System Provenance and Data Collection
@@
System provenance analysis [king:2003sosp, inam2023sok] leverages data collection agents on end-hosts to collect interaction events among key system resources: processes, files, and network sockets. This work relies on in-house data from 21 hosts in a university environment, the DARPA Transparent Computing dataset [darpa], and datasets from previous study [mukherjee2023sec]. Our in-house data collection accumulates 13 GB to 92 GB daily, tracking around 875 unique programs, 7,025K processes, 4,824K network connections, and 111,583K file operations.
@@
Our system provenance data schema, detailed in Table I, is similar to DARPA’s Common Data Model (CDM) [darpa] schema, but we omit memory objects, registry events, and thread distinctions within a process. These choices were made to balance real-world overhead constraints of load balancing and storage. We also established a malicious testbed to generate malware execution traces. To ensure the freshness and realism of our malware samples, we utilize Cyber Threat Intelligence (CTI) feeds [attck], the VirusTotal database [vtotal], and penetration testing tactics, techniques, and procedures (TTPs) [metasploit, cobaltstrike, immunity]. We refer to DARPA TC  dataset [darpa] and previous literature of [mukherjee2023sec] for APTs.
2.2 Provenance-based ML Security Solutions
@@
Recent advancements in system event collection have revitalized host-based IDSs. Although initially proposed in the late ’90s [self1996sp], host-based IDSs have proved effective against advanced attacks such as APTs and Fileless Malware . Various studies [wang2020ndss, sigl2021sec, yang2023prographer, zengy2022shadewatcher, cheng2024kairos, rehman2024flash, goyal2024rcaid] demonstrate the efficacy of provenance-based ML techniques in identifying behavioral deviations. However, the high detection performance is offset by a lack of trustworthy explanations and high false positive rates, impacting operational confidence and resource allocation. Detection tasks are prone to errors and are difficult to verify without full knowledge of typical behaviors. Our work focuses on enhancing the transparency and explainability of provenance-based ML security. We concentrate on a balanced set of security tasks of malware and APT detection. In this paper, we consider graph-level tasks; providing explanations for node and edge level tasks is a promising direction for future work that aligns with upcoming security detection system trends.
@@
GNNs for System Provenance. Our research employs an industry-standard GNN framework (i.e., Deep Graph Learning (DGL)) [dgl] to model and explain system provenance graphs. Leveraging DGL’s mature development ecosystem ProvExplainer aligns with current analytical techniques and streamlined integration into real-world applications. Despite the security community’s historical preference for custom detectors [^1] It has only been recently that GNNs have started gaining attention within the community. [yang2023prographer, zengy2022shadewatcher, cheng2024kairos, rehman2024flash, goyal2024rcaid, jia2024magic] due to the complexity and heterogeneity of provenance graphs, we chose a general GNN framework for universal applicability and broader impact. Notably, the DGL community has integrated our heterogeneous GNN enhancement [dgl-gn-het, dgl-pg-het, dgl-pg-hom, dgl-sub-hom, dgl-sub-het].
2.3 Explainability and Security Applications

### v7

@@ Abstract
Advanced cyber threats (e.g., Fileless Malware and Advanced Persistent Threat (APT)) have driven the adoption of provenance-based security solutions. These solutions employ ML models for behavioral modeling and critical security tasks such as malware and anomaly detection. However, the opacity of ML-based security models limits their broader adoption, as the lack of transparency in their decision-making processes restricts explainability and verifiability. We tailored our solution towards Graph Neural Network (GNN)-based security solutions since recent studies employ GNNs to comprehensively digest system provenance graphs for security critical tasks.
@@
To enhance the explainability of GNN-based security models, we introduce ProvExplainer, a framework offering instance-level security-aware explanations using an interpretable surrogate model. ProvExplainer’s interpretable feature space consists of discriminant subgraph patterns and graph structural features which can be directly mapped to the system provenance problem space, making the explanations human understandable. Considering both the subgraph patterns and graph structural features, gives ProvExplainer the unique advantage of providing explanations that are sensitive to both local and global contexts.
@@
By considering prominent GNN architectures (e.g., GAT and GraphSAGE ) for anomaly detection tasks, we show how ProvExplainer outperformed the current State-Of-The-Art (SOTA) GNN explainers to deliver security domain and instance-specific explanations. We measure the explanation quality using $\text{fidelity}^{+}$ / $\text{fidelity}^{-}$ metric as used by traditional GNN explanation literature, we incorporate the precision/recall metric where we consider the accuracy of the explanation against the ground truth, and we designed a human actionability metric based on graph traversal distance. On real-world Fileless Malware and APT datasets, ProvExplainer achieves up to 29%/27%/25%/1.4x higher $\text{fidelity}^{+}$ , precision, recall and actionability (where higher values are better), and 12% lower $\text{fidelity}^{-}$ (where lower values are better) when compared against SOTA GNN explainers.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2024 26 February - 1 March 2024, San Diego, CA, USA ISBN 1-891562-93-2 https://dx.doi.org/10.14722/ndss.2024.23xxx www.ndss-symposium.org
I Introduction
Given the significant threat posed by advanced and sophisticated adversaries [1, 2, 3, 4] (i.e., malware writers and APT actors), security is paramount in today’s highly digitized society. Recent advances in system monitoring have generated a variety of fine-grained telemetry for advanced security analysis. To leverage this telemetry data to counter threats, various learning-based security tools [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] have been deployed, with provenance-based solutions being the most notable.
@@ I Introduction
System provenance captures causally dependent system events on an host, represented as a provenance graph. Provenance graphs are heterogeneous, comprising various node types (e.g., process, file, and socket) and edge types (e.g., read, write, and send), each with textual and numerical attributes, effectively representing dynamic runtime behaviors. Hence, provenance graphs provide irreplaceable security defense for countering sophisticated and stealthy APT attack campaigns. Particularly, Endpoint Detection and Response (EDR) solutions [22, 23, 24], built on fine-grained system provenance data, have become a mainstream security defense for enterprises.
@@
Recent advancements in GNNs enable the direct learning of relational dependencies and topological structures from graphs. GNNs utilize a permutation-invariant message-passing mechanism to aggregate information from node neighborhoods, expanding the information propagated through each neighborhood with every iteration. This allows GNNs to learn from complex graph structures automatically, which has proven beneficial for provenance solutions [21, 20, 19, 25, 26, 27].
While GNN-based security models are highly effective, their opaque nature significantly limits their advantages, i.e., due to the lack of clear and verifiable explanations. These models fail to provide ground truth relevant explanations, which are essential for building trust among security practitioners. In this context, ground truth relevant explanations refer to explanations aligned with specific system actions that represent Tactics, techniques, and Proceduress (TTPs), and system artifacts that are familiar to analyst and can be validated by existing security intelligence (e.g., attack memos [28, 29] and VirusTotal [30] reports).
While studies [31, 32, 33] have explored domain-agnostic explanations for GNN decisions, they lack validation in the security domain, where contextualizing explanations with program behaviors is critical for trusting the underlying model. Recent research [34, 35, 36] has emphasized verifiable explanations for neural networks in binary analysis and vulnerability discovery, yet our study is the first to validate explanations within the system provenance domain.
Explanations have different meanings for model developers and end-users, since end-users cannot make sense of explanations in terms of model weights, but can make sense if explanations are in terms of edges or subgraphs. Therefore, in the security domain, models must explain their decision-making processes [37] that are verifiable using the ground truth and domain expertise.
@@
In this paper, we introduce ProvExplainer, a novel approach to enhance the explainability of GNN-based detection models trained on system provenance data by utilizing an interpretable surrogate model, specifically Decision Trees (DT), with security-aware discriminant subgraph patterns and graph structural features. We efficiently mined subgraph patterns [38] using a graph evolution miner (GERM [39]) to identify discriminative subgraph patterns between malicious and benign datasets. Our analysis revealed that these subgraph patterns correspond to well-known malicious activities, such as system probing [40], and malware replication [41] and staging [42]. Additionally, we extracted structural graph features, such as degree centrality, betweenness centrality, closeness centrality, eigenvector centrality, clustering coefficient, and clustering triangles. These structural features capture long-term dependencies within the system, while the subgraph patterns represent local dependencies. Together, they provide explanations for both short-term and long-term dependencies.
ProvExplainer is specifically designed to address the unique challenges posed by provenance graphs—complex, heterogeneous, and large-scale structures that capture dynamic system behaviors with fine-grained causal semantics. Unlike prior efforts that apply explanation techniques to general-purpose graph domains, our work introduces a domain-specialized explanation framework tailored for provenance-based GNN intrusion detection. Provenance graphs differ significantly from other graph modalities due to their temporal evolution, multi-typed nodes and edges, and security-relevant semantics embedded in system events. These aspects introduce non-trivial complexities, such as the need to capture temporal causality, reason over multi-hop dependencies, and interpret system-level semantics in a trustworthy manner.
ProvExplainer addresses these challenges by adapting DT-based surrogate models with domain informed features, discriminative subgraph patterns and graph-theoretic measures, that are aligned with operational security workflows. For security analysts, ProvExplainer offers three critical benefits: (1) it builds confidence that the GNN bases its predictions on legitimate anomalous subgraphs in the provenance graph; (2) it enables analysts to investigate alerts by localizing explanations to meaningful subgraphs representative; and (3) it equips operators with interpretable justifications to support triage and response decisions.
@@
We emphasize that ProvExplainer is an explanation framework, not an attack detection technique. Unlike GNNs, surrogate Decision Trees (DTs) sacrifice generalizability, expressiveness, and automatic feature extraction to gain interpretability, making them unsuitable for direct attack detection. ProvExplainer is a GNN-structure-agnostic framework capable of incorporating an arbitrary number of subgraph patterns and structural graph features, and it is designed for easy extensibility with additional security-aware graph features.
We focus on the structural features of system provenance graphs, without considering the text or numerical node and edge attributes. We restrict the scope of our attribute view for the following reasons: first, the core merit of provenance-based Intrusion Detection System (IDS) lies in the ability to retain features drawn from causal dependencies captured in graph structural relations; these structural features are significantly more difficult for adversaries to manipulate than textual attributes (e.g., filenames). Secondly, if we cannot explain GNN decisions based solely on graph structures, it becomes even more challenging to interpret decisions when GNNs incorporate both structural and attribute-based complexities.
We evaluated the effectiveness of ProvExplainer using $\text{fidelity}^{+}$ $\text{fidelity}^{-}$ , following prior GNN explainability research [43]. An effective explainer should have a higher $\text{fidelity}^{+}$ , as it quantifies the prediction change when important subgraphs are removed, and a lower $\text{fidelity}^{-}$ , as it reflects the prediction change when non-important subgraphs are excluded. We also incorporated two additional measures, precision and recall, that captures the accuracy of the explanation provided compared against the ground truth. Finally, we introduce a human actionability metric that measures the graph traversal distance between an explanation and the ground truth artifacts; a lower score indicates that the explanation guides analysts more directly to the relevant evidence, thereby reducing their investigative workload. In our experimental evaluation, we show ProvExplainer surpasses current SOTA GNN explainers in the majority of our security-oriented evaluation tasks.
@@
To demonstrate the advantages of ProvExplainer in the system provenance domain, we compared ProvExplainer against SOTA GNN explainers such as GNNExplainer [31], PGExplainer [32], and SubgraphX [33] across well-balanced security domain tasks, including malware detection, and APT detection. These SOTA GNN explainers provide the best performance according to other security studies [44, 35]. The GNN models in our evaluation are trained with an extensive dataset collected using our in-house deployment which monitored system events on an average of 86 mixed Windows and Linux hosts over 13 months. In addition, in our evaluation, we used the DARPA [darpa] Transparent Computing dataset, the APT dataset from [46], and malware lists from [47].
@@
To the best of our knowledge, our research is the first to leverage security-aware discriminant subgraph pattern and graph structural features specifically tailored for human interpretation of GNN-based detection models. We summarize the contributions of ProvExplainer as follows:
@@
  • ProvExplainer utilizes discriminant subgraph patterns and graph structural features to provide explanation for GNN-based detection models.
@@
  • Our security-aware features enabled surrogate DTs to achieve 95% agreement with the GNN detection model.
  • ProvExplainer increased $\text{fidelity}^{+}$ (e.g., higher is better) by 29%, precision by 27%, recall by 25%, human actionability by 1.4x and decreased $\text{fidelity}^{-}$ (e.g., lower is better) by 12% compared to SOTA GNN explainers.
@@
  • We curated an extensive dataset of APT scenarios ( i.e., DARPA [darpa] and past literature [46]) and collected real-world malware samples from [47] to confirm the generalizability of ProvExplainer.
II Background and Related Work
II-A System Provenance and Data Collection
@@
System provenance analysis [king:2003sosp, 49] leverages data collection agents on end-hosts to collect interaction events among key system resources: processes, files, and network sockets. This work relies on in-house data from 86 hosts in a university environment, the DARPA Transparent Computing dataset [darpa], and datasets from previous study [46]. Our in-house data collection accumulates 13 GB to 92 GB daily, tracking around 875 unique programs, 7,025K processes, 4,824K network connections, and 111,583K file operations.
@@
Our system provenance data schema, detailed in Table III, is similar to DARPA’s Common Data Model (CDM) [darpa] schema, but we omit memory objects, registry events, and thread distinctions within a process. These choices were made to balance real-world overhead constraints of load balancing and storage. We also established a malicious testbed to generate malware execution traces. To ensure the freshness and realism of our malware samples, we utilize Cyber Threat Intelligence (CTI) feeds [attck], the VirusTotal database [vtotal], and penetration testing TTP [metasploit, cobaltstrike, immunity]. We refer to DARPA TC dataset [darpa] and previous literature of [46] for APTs.
II-B Provenance-based ML Security Solutions
@@
Recent advancements in system event collection have revitalized host-based IDSs. Although initially proposed in the late ’90s [54], host-based IDSs have proved effective against advanced attacks such as APTs and Fileless Malware . Various studies [15, 18, 55, 56, 19, 20, 21, 26, 27, 57] demonstrate the efficacy of provenance-based Machine Learning (ML) techniques in identifying behavioral deviations. However, the high detection performance is offset by a lack of trustworthy explanations and high false positive rates, impacting operational confidence and resource allocation. Our work focuses on enhancing the transparency and explainability of provenance-based ML security. We concentrate on a balanced set of security tasks of malware and APT detection. In this paper, we consider graph-level tasks; providing explanations for node and edge level tasks is a promising direction for future work that aligns with upcoming IDS trends.
@@
GNNs for System Provenance. Our research employs an industry-standard GNN framework (i.e., Deep Graph Learning (DGL)) [58] to model and explain system provenance graphs. Leveraging DGL’s mature development ecosystem ProvExplainer aligns with current analytical techniques and streamlined integration into real-world applications. Despite the security community’s historical preference for custom detectors [55, 56, 19, 20, 21, 25, 27, 26] due to the complexity and heterogeneity of provenance graphs, we chose a general GNN framework for universal applicability and broader impact. Notably, the DGL community has integrated our heterogeneous GNN enhancement.
II-C Explainability and Security Applications